### v1

# SALT: Parameter-Efficient Fine-Tuning via Singular Value Adaptation with Low-Rank Transformation

Abstract
The complex nature of medical image segmentation calls for models that are specifically designed to capture detailed, domain-specific features. Large foundation models offer considerable flexibility, yet the cost of fine-tuning these models remains a significant barrier. Parameter-Efficient Fine-Tuning (PEFT) methods, such as Low-Rank Adaptation (LoRA), efficiently update model weights with low-rank matrices but may suffer from underfitting when the chosen rank is insufficient to capture domain-specific nuances. Conversely, full-rank Singular Value Decomposition (SVD) based methods provide comprehensive updates by modifying all singular values, yet they often lack flexibility and exhibit variable performance across datasets. We propose SALT (Singular Value Adaptation with Low-Rank Transformation), a method that selectively adapts the most influential singular values using trainable scale and shift parameters while complementing this with a low-rank update for the remaining subspace. This hybrid approach harnesses the advantages of both LoRA and SVD, enabling effective adaptation without relying on increasing model size or depth. Evaluated on 5 challenging medical datasets, ranging from as few as 20 samples to 1000, SALT outperforms state-of-the-art PEFT (LoRA and SVD) by 2% to 5% in Dice with only 3.9% trainable parameters, demonstrating robust adaptation even in low-resource settings. The code for SALT is available at: https://github.com/BioMedIA-MBZUAI/SALT.
Keywords:
Singular Value Decomposition, Parameter-Efficient Fine-Tuning, LoRA, Foundation Models, Domain Adaptation.
1 Introduction
Medical image segmentation plays a pivotal role in modern healthcare by enabling precise delineation of anatomical structures, pathological regions, and functional tissues across diverse imaging modalities such as CT, MRI, X-ray, ultrasound, and endoscopy [11, 20]. While Deep Learning (DL) has revolutionized this field, with architectures like U-Net [20] and nnU-Net [8] achieving state-of-the-art (SOTA) performance through sophisticated designs and extensive data augmentation, these models often rely on large parameter counts, limiting their adaptability to new datasets or modalities. For instance, nnU-Net’s success hinges on dynamic architecture adjustments and aggressive augmentation pipelines, while DeepLabv3+ [3] employs computationally intensive components like spatial pyramid pooling. Recent advances in foundation models have introduced specialized tools for diverse computer vision tasks. CLIP [18] and ALIGN [9], trained on large-scale image-text datasets, excel at cross-modal understanding for zero-shot image classification and open-set retrieval. In contrast, the Segment Anything Model (SAM) [10, 19], trained on billions of natural images, exhibits robust zero-shot generalization capabilities and initially raised hopes for universal segmentation. However, SAM’s performance on medical data has proven suboptimal, with studies [26, 6] demonstrating its inferiority to specialized models like U-Net++ [27], particularly in Dice scores across 12 medical datasets. Even with prompt engineering, SAM’s zero-shot accuracy remains below fully supervised methods [4], underscoring the domain gap between natural and medical images.
Recent efforts to adapt SAM to medical imaging, such as MedSAM [13], SAM Adapter [23], and S-SAM [16], have made strides by incorporating medical training data, low-rank adapters, or SVD-based tuning. Yet challenges persist: PEFT methods must balance minimal parameter overhead, computational feasibility, and preservation of pre-trained knowledge while addressing unique medical imaging challenges like speckle noise, low contrast, and modality-specific artifacts. Promising approaches that arose to adapt foundational models, especially language models, like LoRA [7] and AdaLoRA [25] approximate weight updates via low-rank matrices but struggle to capture both dominant and nuanced features, while SVD-based methods like PiSSA [15] focus on principal singular vectors to capture the most significant data patterns, but this rigid prioritization can limit their adaptability to dynamic or fine-grained data characteristics. To address these limitations, we propose Singular Value Adaptation with Low-Rank Transformation (SALT), a novel PEFT framework that synergizes the strengths of SVD and low-rank adaptation. SALT selectively scales critical singular values to preserve foundational knowledge while applying trainable low-rank transformations to residual components, enabling efficient adaptation to medical domains with minimal parameter overhead. Evaluated on five challenging medical datasets, with sample sizes ranging from 20 to 1000 and demanding a high level of detailed segmentation, SALT consistently outperforms LoRA and SVD, achieving a 2 to 5% improvement in Dice score while utilizing only 3.9% trainable parameters. Our contributions are threefold:
  • We introduce SALT, a hybrid PEFT method utilizing SVD-based dominant singular value adaptation with low-rank residual updates, achieving parameter-efficient domain adaptation.
  • We demonstrate SALT’s utility through extensive experiments on diverse medical datasets from various modalities, showing consistent improvements over state-of-the-art PEFT methods.
  • We present a comprehensive study exploring how singular value distributions and trainable parameters’ allocation impact domain-specific fine-tuning for different challenging medical datasets.
2 Methodology
Figure 1: Overview of the SALT architecture. Starting from the segmentation model and providing a detailed view of the SALT mechanism within the Multi-Head Attention layers of the image encoder.
Problem Formulation. Let S denote SAM model with pre-trained parameters $\Theta$, which consists of an image encoder ($\mathcal{E}_{\mathrm{i}}$), a prompt encoder ($\mathcal{E}_{\mathrm{p}}$), a mask decoder ($\mathcal{D}_{\mathrm{m}}$). $\mathcal{E}_{\mathrm{i}}$ contains $L$ blocks, each comprising a multi-head attention module and a Multi-Layer Perceptron (MLP). For a given input image $\mathbf{i}$ and prompt embedding $\mathbf{p}$, the encoder $\mathcal{E}_{\mathrm{i}}$ produces intermediate feature maps $\{\mathbf{f}^{(l)}\}_{l=1}^{L}$. Concretely, in the $l$-th block, the main operations can be written as:
$q,k,v=W_{qkv}^{l}\,i+b_{qkv}^{l},\quad o_{n}=W_{M}^{l}\,i+b_{M}^{l},$ (1)
where $W_{qkv}^{l},W_{M}^{l}\in\mathbb{R}^{D\times K}$ and $b_{qkv}^{l},b_{M}^{l}\in\mathbb{R}^{D}$ are the attention/MLP weights and biases for block $l\in\{1,...,L\}$. While effective for natural images, the pre-trained weights $W\in\mathbb{R}^{D\times K}$ may not capture all domain-specific nuances.
2.0.1 SALT.
Through SALT, we refine these weights using the following procedure. First, consider a weight matrix $\mathbf{W}^{(l)}$. SALT decomposes $\mathbf{W}^{(l)}$ using Singular Value Decomposition (SVD).
$\mathbf{W}^{(l)}=\mathbf{U}^{(l)}\,\mathbf{\Sigma}^{(l)}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top},$ (2)
where $\mathbf{U}^{(l)}\in\mathbb{R}^{D\times D}$ (left singular vectors), $(\mathbf{V}^{(l)})^{\top}\in\mathbb{R}^{K\times K}$ (right singular vectors), and $\mathbf{\Sigma}^{(l)}\in\mathbb{R}^{D\times K}$ is a diagonal matrix of the singular values corresponding to each singular vector. We denote $min(D,K)$ as $r_{max}$ and split this matrix into two matrices (i) $\mathbf{\Sigma}^{(l)}_{r}$ which contains a diagonal vector of top r singular values where $r\leq r_{max}$ (ii) $\mathbf{\Sigma}^{(l)}_{r^{\prime}}$ containing the remaining $r^{\prime}$ singular values where $r^{\prime}=r_{max}-r$. Then SALT applies (i) trainable scale and shift transformations to the top $r$ singular values. Concretely, define parameters $\alpha^{(l)}$ and $\beta^{(l)}$ as a diagonal matrices $\in\mathbb{R}^{r\times r}$ to transform $\mathbf{\Sigma}^{(l)}_{r}$:
$\mathbf{\Sigma}^{{}^{\prime}(l)}_{r}=\alpha^{(l)}\odot\mathbf{\Sigma}^{(l)}_{r}+\beta^{(l)},$ (3)
where $\odot$ denotes Hadamard product, and (ii) a LoRA-based low-rank update $\mathbf{X}^{(l)}\mathbf{Y}^{(l)\top}$ where $\mathbf{X}^{(l)}\in\mathbb{R}^{r^{\prime}\times d_{\text{lora}}}$ , $\mathbf{Y}^{(l)}\in\mathbb{R}^{d_{\text{lora}}\times r^{\prime}}$ to the remaining singular values $\mathbf{\Sigma}^{(l)}_{r^{\prime}}$. The updated modified singular matrix $\mathbf{\Sigma}^{{}^{\prime}(l)}$ is then activated elementwise by $\mathrm{ReLU}$ to preserve semi-definiteness, yielding the updated weight matrix:
Figure 2: Comparison of PEFT methods: (A) Full fine-tuning updates all $W$ parameters. (B) LoRA updates low-rank matrices. (C) SVD fine-tunes scale and shift components of $W$. (D) SALT (Ours) integrates scale, shift, and low-rank updates for singular values.
$\widetilde{\mathbf{W}}^{(l)}\;=\;\mathbf{U}^{(l)}\,\bigl{[}\mathrm{ReLU}\!\bigl{(}{\mathbf{\Sigma}}^{{}^{\prime}(l)}\bigr{)}\bigr{]}\,\bigl{(}\mathbf{V}^{(l)}\bigr{)}^{\top}.$ (4)
During fine-tuning, only the newly introduced SALT parameters $\{\alpha^{(l)},\beta^{(l)},\mathbf{X}^{(l)}$, $\mathbf{Y}^{(l)}\}$ are optimized along with Normalization Layers and Text Affine Layers, while all original SAM weights remain frozen. The prompt encoder $\mathcal{E}_{\mathrm{p}}$, receiving text prompts $\mathbf{p}$, yields additional embeddings merged with the image-encoder $\mathcal{E}_{\mathrm{i}}$ output in the mask decoder $\mathcal{D}_{\mathrm{m}}$. We define SALT, denoted $S^{\prime}$, as the combination of two components: The original SAM model’s frozen output $S(i;\Theta_{\text{fixed}})$, which remains static during fine-tuning, and a lightweight, learnable adjustment term $\Delta S(i;\theta)$, introduced by SALT layers,capturing task-specific refinements through the trainable parameters $\theta$. This formulation ensures that the core SAM backbone remains unchanged, while task-specific adaptations are layered on top via the SALT modules. Putting this together, our objective is to minimize the loss function $\mathcal{L}(S^{\prime}(i),y)$ where $i$ is an input sample and $y$ is the corresponding ground truth label. To train SALT, we use a combination of segmentation losses, including a Focal loss, a Dice loss, and a regularization term expressed as:
$\mathcal{L}_{\text{reg}}=\underbrace{\|\alpha\odot s+\beta-s\|_{F}}_{\text{Scale-Shift}}+\underbrace{\|XY\|_{F}}_{\text{LoRA}},$ (5)
where $\|\cdot\|_{F}$ denotes Frobenius norm. This dual constraint balances two goals: retaining the useful, pre-trained features and making only the minimal modifications needed for effective domain adaptation.
Rationale for the SALT Design As shown in Fig. 2, LoRA [7] modifies a weight matrix $W\in\mathbb{R}^{D\times K}$ by adding $XY$, where $X\in\mathbb{R}^{D\times r}$ and $Y\in\mathbb{R}^{r\times K}$ have rank $r\ll\min(D,K)$. This reduces trainable parameters from $D\times K$, down to $r(D+K)$, but can underfit if $r$ is too small. By contrast, SVD-based fine tuning in S-SAM [16], shown in Fig. 2, learns only the scale and shift of each singular value in $W$. Since $\mathrm{rank}(W)=\min(D,K)$, this corresponds to $2\times\min(D,K)$ parameters, fewer than LoRA, yet fully updating every singular value may lack flexibility for subtle domain shifts. SALT in Fig. 2 merges the strengths of both. It trains the top singular values via scale and shift and uses LoRA blocks for the lower singular subspace. This dual strategy retains key features while capturing domain-specific nuances, balancing parameter efficiency with adaptation capability.
3 Experiments
Datasets. In this work, we use datasets spanning three anatomical domains: neurovascular analysis with DIAS [12], a dynamic 2D+time DSA dataset for intracranial artery segmentation in cerebrovascular disease (20 train/10 test). Retinal assessment via ROSE [14] for OCT-Angiography-based microvasculature analysis (22 train/8 test). DRIVE [22] for diabetic retinopathy screening in RGB fundus images (14 train/6 test). Cardiovascular studies using ARCADE [17] (700 train/300 test) and XRay-Angio [2] (93 train/41 test), both leveraging X-ray angiography, where the former is for region-based coronary artery disease diagnostics and the latter is for multiscale segmentation of occluded vessels. The datasets were randomly partitioned into training and testing subsets to ensure unbiased evaluation.
Implementation Details. We apply SALT to the image encoder layers of the prompt-adapted SAM with a 512 input size and text-based prompts. Following the pipeline in S-SAM [16], training uses AdamW ($\text{LR}=1\times 10^{-4}$, weight decay $=1\times 10^{-2}$), batch size 5, and 200 epochs, with a step decay scheduler for learning rate adjustment. Data augmentation includes resizing, random flips, ±$\pm 10^{\circ}$ rotations, and brightness adjustments. The total loss combines focal loss, Dice loss, and regularization loss (scaled by 0.01). Evaluation metrics include the Dice Similarity Coefficient (DSC) and 95th percentile Hausdorff Distance (HD95). Experiments were conducted on an NVIDIA RTX 4090 (24GB).
4 Results and Discussion
Table 1: Performance comparison of traditional DL and SAM-based PEFT methods using Dice and HD95 metrics on a variety of medical segmentation datasets.
Model Rank %Trainable ARCADE DIAS ROSE XRay-Angio Drive Average
Dice $\uparrow$ HD95 $\downarrow$ Dice $\uparrow$ HD95 $\downarrow$ Dice $\uparrow$ HD95 $\downarrow$ Dice $\uparrow$ HD95 $\downarrow$ Dice $\uparrow$ HD95 $\downarrow$ Dice $\uparrow$ HD95 $\downarrow$
Traditional DL Methods
U-Net [20] 100% / 6.5M 0.36 157.14 0.13 262.66 0.63 8.07 0.71 30.40 0.18 84.43 0.40 108.54
UNETR [5] 100% / 91M 0.48 162.44 0.31 183.97 0.63 8.96 0.74 20.62 0.46 33.58 0.52 81.91
RegUnet [20] 100% / 7.8M 0.11 247.79 0.35 172.09 0.22 58.08 0.42 79.51 0.43 72.97 0.31 126.08
SegNet [1] 100% / 29M 0.14 231.11 0.13 295.06 0.38 9.09 0.09 110.94 0.14 79.03 0.18 145.05
DeepLabV3+ [3] 100% / 45M 0.55 103.22 0.55 77.66 0.52 12.90 0.74 14.03 0.46 27.79 0.56 47.12
Segformer [24] 100% / 43M 0.53 113.98 0.57 73.64 0.56 11.21 0.75 16.55 0.48 36.98 0.58 50.47
SAM-Based PEFT Methods (241M)
SAM w/ text prompts 0.00 - 0.00 - 0.00 - 0.04 - 0.00 - 0.01 -
S-SAM [16] 0.40% / 1M 0.76 52.44 0.67 42.36 0.65 12.41 0.75 27.56 0.70 15.81 0.71 30.12
LoRA [7] 4 0.64% / 1.5M 0.78 39.58 0.61 27.81 0.59 16.26 0.73 26.78 0.62 21.04 0.66 26.29
LoRA [7] 256 14.08% / 33.9M 0.81 36.89 0.63 34.36 0.62 14.62 0.76 24.84 0.68 19.00 0.70 25.94
SALT (ours) 4 0.46% / 1.1M 0.78 51.74 0.69 39.58 0.66 12.19 0.75 29.55 0.72 16.04 0.72 29.82
SALT (ours) 256 3.9% / 9.4M 0.81 42.50 0.71 27.35 0.67 12.64 0.77 23.67 0.75 13.20 0.74 23.87
We evaluate the performance of SALT against SOTA PEFT techniques, specifically LoRA [7] and S-SAM [16]. As shown in Table 1, our experimental evaluation demonstrates the effectiveness of SAM-based PEFT methods compared to traditional deep learning approaches across five medical imaging datasets. While traditional methods like DeepLabV3+ [3] and U-Net [20] showed competitive performance on datasets such as XRay-Angio [2], they require training of all parameters. On the other hand, SALT achieves superior performance with significantly fewer trainable parameters. With rank-256 configuration, attaining an average Dice score of 0.74 while requiring only 3.9% of parameters to be trained. This represents a substantial improvement over both traditional methods and other parameter-efficient approaches like LoRA [7] and S-SAM [16]. Notably, SALT consistently achieves higher Dice scores over other methods across all datasets, achieving the highest scores on DIAS (0.71), ROSE (0.67), XRay-Angio (0.77), and Drive (0.75). The rank configuration comparison reveals that while rank-4 SALT delivers competitive results with minimal parameter overhead (0.46% trainable), the rank-256 version provides optimal performance with a modest increase in trainable parameters. Our analysis of boundary accuracy via HD95 further underscores SALT’s performance. While traditional methods exhibited high HD95 errors, such as SegNet [1] scoring 145.05 and RegUnet [20] scoring 126.08, SAM-based PEFT approaches showed improvements, with SALT achieving the lowest average HD95 ($23.87$) against LoRA[7] ($25.94$) and S-SAM [16] ($30.12$). Notably, SALT demonstrated Strong boundary adherence in challenging cases. SALT’s 13.20 HD95 on DRIVE [22] against S-SAM [16] achieving 15.81 reflects precise vessel delineation in low-contrast retinal images. This dual advantage in both Dice and HD95 confirms SALT’s ability to maintain structural accuracy and boundary precision while operating in parameter-efficient regimes.
Qualitative Analysis. As illustrated in Fig 3, the qualitative results highlight the SOTA performance of our proposed SALT method across various medical imaging datasets. Compared to traditional approaches like UNETR [5] and U-Net [20], as well as other SAM-based PEFT methods such as S-SAM [16] and LoRA [7], SALT consistently delivers more precise vessel segmentation. In the ROSE [14] dataset, SALT excels at capturing the intricate, web-like patterns of retinal vessels while maintaining structural continuity with fewer false positives. For the DIAS [12] dataset, SALT effectively segments complex branching patterns, preserving both main vessel trunks and subtle secondary branches. In X-ray angiography [2] images, SALT achieves clear vessel boundaries with fewer false positives compared to traditional methods. Additionally, in the Drive [22] dataset, it successfully segments retinal vessels despite varying widths and complex backgrounds.
Figure 3: Qualitative Vessel Segmentation. The figure displays segmentation results for five datasets (Arcade [17], Rose [14], Dias [12], X-ray Angio [2], Drive [22]). Columns show the ground truth, SALT, SAM-based PEFT methods, and traditional DL models. Predicted masks are white, and false negatives are red.
Figure 4: Ablation study on SALT configurations And adaptability to SAM2. a) Evaluate model performance and efficiency across LoRA ranks, with the x-axis for parameter combinations, the y-axis for Dice scores, and circle size for trainable parameter percentage. The numbers on the x-axis represent the rank of the top singular values, which are modified using scale and shift in each mentioned layer Conv2d, MLP, and MHA. b) Compares SALT, LoRA [7], and SVD PEFTs integrated with SAM2 on two datasets of different sizes.
Extension to SAM2 [19]. We extend SALT to SAM2 and perform ablation studies comparing LoRA [7], SVD-based fine tuning, and SALT on the Hiera [21] image encoder. The SAM2 [19] image encoder utilizes a hierarchical architecture, capturing multi-scale features from input images with variable embedding dimensions ranging from low-dimensional, early-stage representations to high-dimensional, later-stage ones. This design enables LoRA to adapt effectively. In contrast, the original SAM [10] image encoder employs a backbone with constant, high-dimensional embeddings, where LoRA exhibited instability and was sometimes outperformed by S-SAM’s SVD-based fine tuning method [16]. In SAM2, however, LoRA consistently outperforms relative to SVD, while SALT achieves competitive performance to LoRA with 2.4 times fewer parameters at rank 256 and 1.8× fewer at rank 4.
Ablation Study. We conducted comprehensive ablation experiments on the Drive dataset [22] to evaluate the impact of different architectural components and rank configurations in SALT. Fig 4(a) presents the results across various combinations of LoRA ranks (4, 64, 256) and SALT ranks for Multi-Head Attention (MHA), neck convolutional layers (NECK-Conv2d), and Multi-Layer Perceptron (MLP) modules. The SALT rank for a given layer refers to the number of dominant singular values in the layer’s weight matrix that are adaptively scaled and shifted during fine-tuning. Key findings are: 1) increasing the LoRA rank from 4 to 256 consistently improved Dice scores ($0.72\ \rightarrow 0.76$) while maintaining a modest parameter increase from 0.46% to 5.86% and 2) reducing MHA and MLP ranks from 700 to 200 further enhanced performance, increasing Dice scores from 0.75 to 0.76 with LoRA rank 256. These results suggest that higher LoRA ranks improve expressiveness, while component-specific rank selection is crucial for optimal performance. The optimal configuration (LoRA rank 256 and components rank 200) achieved a Dice score of 0.76 with only 5.84% of trainable parameters, demonstrating SALT’s effective balance of performance and efficiency.
5 Conclusion
In this paper, we introduce SALT, a novel PEFT framework that adapts the Segment Anything Model (SAM) to domain-specific tasks, particularly medical image segmentation. By utilizing SVD-based scaling and shifting of the top singular values with LoRA-based low-rank updates, SALT achieves efficient and flexible adaptation, enabling precise segmentation with minimal computational overhead. Experiments show that SALT outperforms state-of-the-art SAM PEFT adaptations while preserving parameter efficiency and adaptability. Future work will extend SALT to other foundation models and explore dynamic rank selection to enable optimal parameter allocation.
6 Acknowledgments
The authors gratefully acknowledges Juan Lugo for his contributions through out the project.